OpenStudy (anonymous):

Explain how to factor the following trinomials forms: x2 + bx + c and ax2 + bx + c. Is there more than one way to factor this? Show your answer using both words and mathematical notation

OpenStudy (anonymous):

In order to factor x^2+bx+c. You will have to find factors of c that add up to b. For example to factor x^2+7x+10 you find factors of 10 that add up to 7. It's easy to see that they are 5 and 2. Therefore x^2+7x+10=(x+5)(x+2) To factor ax^2+bx+c, you need to find factors of ac that add up to b. But you now have to separate the b terms into a sum of terms using the factors that you have found. To illustrate, let's take 6x^2+11x+3. ac=6*3=18. 2 factors of 18 that add up to 11 are 9 and 2. Now we have 6x^2+11x+3=6x^2+9x+2x+3=3x(2x+3)+(2x+3)=(3x+1)(2x+3).

OpenStudy (anonymous):

how about monomial ?